OpenStudy (anonymous):

Does 4x^2-25 = (2x-5)^2?

OpenStudy (anonymous):

no

OpenStudy (angela210793):

no cause (2x-5)^2=4x^2-20x+25

OpenStudy (anonymous):

angela's explanation is correct

OpenStudy (anonymous):

heck no. left hand side is difference of two squares, right hand side is a perfect square

OpenStudy (anonymous):

4x^2-25=(2x-5)^2 4x^2-25=4x^2-20x+25 4x^2-25-4x^2+20x-25=0 20x-50=0 x=50/20 x=5/2

OpenStudy (anonymous):

@aditya he just asked the question whether the LHS was equal to RHS or not.... this is not to be solved as u hv done...